Nielsen Associates provides marketing services for a number of small manufacturing firms. Nielsen receives a commission of 10 percent of sales. Operating costs are as follows:

Unit-level costs $0.02 per sales dollar
Sales-level costs $100 per sales order
Customer-level costs $800 per customer per year
Facility-level costs $60,000 per year

(a) Determine the minimum order size in sales dollars for Nielsen to break even on an order.

(b) Assuming an average customer places four orders per year, determine the minimum annual sales required to break even on a customer.

(c) What is the average order size in (b)?

(d) Assuming Nielsen currently serves 100 customers, with each placing an average of four orders per year, determine the minimum annual sales required to break even.

(e) What is the average order size in (d)?

(f) Choose the correct answer: "Explain the differences in the answers to (a), (c), and (e)?"

- In multiple customer firms the break-even point decreases as the number of customers increases

- The most important costs to cover are unit level costs

- In the long-run the most important costs are facility level costs

- Even if individual orders have a positive contribution, some customers may be unprofitable
Answer #1

ANSWER

a The minimum order size in sales dollars for Nielsen to break even on an order = sale level cost per order / (commission per unit-unit level cost per unit)
= 100/(0.1-0.02)
1250
b The minimum annual sales required to break even on a customer = customer level cost /(commission per unit-unit level cost per unit)
=((100*5)+800)/(0.1-0.02)
16250
c Average Order size
=16250/5
3250
d The minimum annual sales required to break even = Total costs / (commission per unit-unit level cost per unit)
=(100*5*100+800*100+60000)/(0.1-0.02)
2375000
e Average Order size
=2375000/(5*100)
4750